OpenStudy (anonymous):

(2.6*10^-3)(4.8*10^-4)/1.3*10^-12

OpenStudy (dumbsearch2):

Luls, no direct answers on OpenStudy @kc_kennylau read the Code of Conduct.

OpenStudy (kc_kennylau):

lol?!

OpenStudy (dumbsearch2):

Maybe she needs help about *how* to do it.

OpenStudy (kc_kennylau):

\[\frac{(2.6*10^{-3})(4.8*10^{-4})}{1.3*10^{-12}}\]\[=\frac{2.6\times4.8}{1.3}\times\frac{10^{-3}10^{-4}}{10^{-12}}\]Then apply the laws of indices

OpenStudy (anonymous):

Ok